てnel Mass pa r ev with KHP Mass Pa re Mas Kep added Initial buve 0-30649 endine 5 Naott se lanty tan Sus peete Cr Mais k P added

please help. how do I find the rest of them thanks!
Answer #1

Ans.) The Chemical Reaction taking place is :

KHC8H4O4 + NaOH ----> KNaC8H4O4 + H2O

Ionic equation :

H+ + OH- -------> H2O

From the reaction we find that KHP reacts with NaOH in 1:1 ratio on the basis of stochiometry.

Hence , 1 mol of KHP = 1 mol of NaOH

Trial 1. Given that,

Mass of KHP = 0.3054 g ,

Volume of NaOH used = (final burette reading - initial burette reading) = 20.9 mL - 5.1 mL = 15.8 mL

Number of moles = Given mass (g) / Molar mass of compound (g/mol)

Thus , we have,

Number of moles of KHP = 0.3054 / 204.22 = 0.001495 = 1.495 x 10-3 moles

Hence ,

1.495 x10-3 moles of KHP x 1 mol of KHP = 1.495 x10-3 moles of NaOH x 1 mol of NaOH

Molarity = Number of moles / Volume used (in L)  

Volume of NaOH used = 15.8 mL = 0.0158 L

Molarity of NaOH = 1.495 x10-3 / 0.0158  = 0.0949 M

Trial 2.

Mass of KHP = 0.3073 g ,

Volume of NaOH used = 22.4 mL - 6.9 mL = 15.5 mL = 0.0155 L

No. of moles KHP = 0.3073 / 204.22 = 0.00151 moles = 1.51 x 10-3 moles

No. of moles KHP = No. of moles NaOH = 1.51 x 10-3 moles

Molarity of NaOH = 1.51 x 10-3 / 0.0155 = 0.0974 M

Trial 3.

Mass of KHP = 0.3032 g ,

Volume of NaOH used = 20.6 mL - 5.5 mL = 15.1 mL = 0.0151 L

No. of moles KHP = 0.3032 / 204.22 = 0.00148 moles = 1.48 x 10-3 moles

No. of moles KHP = No. of moles NaOH = 1.48 x 10-3 moles

Molarity of NaOH = 1.48 x 10-3 / 0.0151 = 0.0980 M

Trial 4.

Mass of KHP = 0.3064 g ,

Volume of NaOH used = 18.0 mL - 2.6 mL = 15.4 mL = 0.0154 L

No. of moles KHP = 0.3064 / 204.22 = 0.00151 moles = 1.50 x 10-3 moles

No. of moles KHP = No. of moles NaOH = 1.50 x 10-3 moles

Molarity of NaOH = 1.50 x 10-3 / 0.0154 = 0.0974 M

Average Molarity of NaOH = (0.0949 + 0.0980 + 0.0980 + 0.0974) / 4 = 0.0969 M \approx 0.097 M

Standard Deviation is given as

Where M bar = average molarity, Mi = individual Molaruty measurements, and N = total number of molarities measured.

Standard Deviation \approx 0.00138